# UNIMO-G: Unified Image Generation through Multimodal Conditional Diffusion

### v1

@@ 6 Limitations
Appendix A Implementation Details
@@
    Initial Training with CC3M Dataset: The CC3M dataset, consists of about 3.3M image-description pairs, was translated into Chinese using Baidu Translate API ***https://fanyi.baidu.com. The model is trained from scratch at 256x256 using the AdamW optimizer with a weight decay of 0.01, a learning rate of 5e-5, and a batch size of 40x256 for 100K steps.
@@
    Expansion with Large-Scale Chinese Data: We incorporate about 300M internal Chinese text-image pairs from multiple datasets, including LAION-2B Schuhmann et al. (2022), COYO-700M Byeon et al. (2022), Conceptual Captions Changpinyo et al. (2021) and a series of internal Chinese datasets. The English captions are translated into Chinese. This stage, with a constant learning rate of 5e-5, initially training at 256x256 resolution for 500K steps with a batch size of 40x256, then progresses to 512x512 for 200K steps with a batch size of 12x256.
@@
    Refinement with High-Quality Corpus: The final stage focuses on fine-tuning on high-quality corpus, and continues training at 512x512 and 1024x1024 resolutions. Selection criteria include an aesthetic threshold above 6.5 (LAION-Aesthetics V2) and an image-text similarity score over 0.5 (CLIP-base-32), resulting in about 1M pairs. A multi-scale training strategy dividing the image size into 5 buckets [0.5, 0.75, 1.0, 1.5, 2.0] supports various image aspect ratios. The final stage uses a learning rate of 1e-5 for 200K steps with a batch size of 3072.
@@
We developed a multimodal-to-image instruction tuning dataset utilizing the 1M high-quality image-text pairs. The process, depicted in Figure 3, involves: (1) Generation of captions and extraction of entity tokens from captions by the MLLM; (2) Identifying entity detection boxes via Grounding DINO Liu et al. (2023); (3) Segmentation and extraction of regions corresponding to each entity by SAM Kirillov et al. (2023); (4) Randomly substitution of entity tokens with their corresponding image segments. This process effectively transforms the original text-image pairs into multimodal-image pairs, which are then employed for refining multimodal instruction tuning. We maintain a learning rate of 1e-5, training for 200K steps with a batch size of 3072. To preserve text-to-image generation capabilities, 10% of the training uses original textual captions.
@@
Figure 7: Illustration of images in MultiBench.
Appendix B Images and Prompts of MultiBench

### v3

@@ 6 Limitations
Appendix A Multi-modal Large Language Model
The structure of our MLLM is similar to MiniGPT-4 Zhu et al. (2023) and BLIP-2 Li et al. (2023a), which consists of a vision encoder with a pretrained ViT-G and BLIP-2 Q-Former, a single linear projection layer, and a pre-trained 6B Transformer-based LLM. The MLLM totally contains about 7.8B parameters. The encoding process of multimodal prompts is illustrated as in Figure 7. Each image is firstly encoded as 32 visual tokens by the visual encoder, and then linearly transformed into the LLM embedding space. The pre-trained LLM is utilized as the context encoder and decoder. Our MLLM underwent training on web-scale multimodal corpora, comprising text corpora, image-caption pairs, and interleaved data of images and texts.
Figure 7: Illustration of the encoding of multimodal prompts by the MLLM.
Appendix B Implementation Details
@@
    Initial Training with CC3M Dataset: The CC3M dataset, consists of about 3.3M image-description pairs, was translated into Chinese using the Baidu Translation API. The model is trained from scratch at 256x256 using the AdamW optimizer with a weight decay of 0.01, a learning rate of 5e-5, and a batch size of 40x256 for 100K steps, which costs about 300 A100 GPU days.
@@
    Expansion with Large-Scale Chinese Data: We incorporate about 300M Chinese text-image pairs from multiple datasets, including LAION-2B Schuhmann et al. (2022), COYO-700M Byeon et al. (2022), Conceptual Captions Changpinyo et al. (2021) and a series of internal Chinese datasets. The English captions are translated into Chinese. This stage, with a constant learning rate of 5e-5, initially training at 256x256 resolution for 500K steps with a batch size of 40x256, then progresses to 512x512 for 200K steps with a batch size of 12x256, which costs about 3000 GPU days.
@@
    Refinement with High-Quality Corpus: The final stage focuses on fine-tuning on high-quality corpus, and continues training at 512x512 and 1024x1024 resolutions. Selection criteria include an aesthetic threshold above 6.5 (LAION-Aesthetics V2) and an image-text similarity score over 0.5 (CLIP-base-32), resulting in about 1M pairs. A multi-scale training strategy dividing the image size into 5 buckets [0.5, 0.75, 1.0, 1.5, 2.0] supports various image aspect ratios. The final stage uses a learning rate of 1e-5 for 200K steps with a batch size of 3072, which costs about 1000 GPU days.
@@
We developed a multimodal-to-image instruction tuning dataset utilizing the 1M high-quality image-text pairs. The process, depicted in Figure 3, involves: (1) Generation of captions and extraction of entity tokens from captions by the MLLM; (2) Identifying entity detection boxes via Grounding DINO Liu et al. (2023b); (3) Segmentation and extraction of regions corresponding to each entity by SAM Kirillov et al. (2023); (4) Randomly substitution of entity tokens with their corresponding image segments. This process effectively transforms the original text-image pairs into multimodal-image pairs, which are then employed for refining multimodal instruction tuning. We maintain a learning rate of 1e-5, training for 200K steps with a batch size of 3072. To preserve text-to-image generation capabilities, 10% of the training uses original textual captions. This process totally costs about 800 A100 GPU days.
Number of Entities Ratios
0 16%
1 36.8%
2 25.8%
3 12.3%
>=4 9.1%
Table 4: Statistics of the number of subject entities in our constructed multi-modal prompts.
Num of Entities Methods DINO CLIP-I CLIP-T Avg
2 BLIP-Diffusion 0.455 0.675 0.249 0.460
KOSMOS-G 0.465 0.704 0.279 0.483
UNIMO-G 0.485 0.699 0.293 0.492
3 BLIP-Diffusion 0.344 0.608 0.249 0.400
KOSMOS-G 0.350 0.621 0.287 0.419
UNIMO-G 0.364 0.614 0.306 0.428
Table 5: Comparisons of the performance with different number of subject entities in a multi-modal prompt on MultiBench.
Figure 8: Illustration of images in MultiBench.
Appendix C Analysis of Multi-modal Prompts
In our analysis of multi-modal prompts, we meticulously examined the composition of our dataset, finding an average of $42.48$ text tokens and $1.67$ subject entities per prompt. The distribution of entity counts revealed a predominance of images featuring fewer than three subject entities, with a minor proportion exceeding this count. The ratios of different number of entities are shown as in Table 4.
In our experiments, we have assessed the model’s performance using DreamBench for single-entity prompts and MultiBench for multi-entity scenarios, specifically focusing on two and three entities. We separately evaluate the performance on two and three entities as shown in Table 5. The results, as detailed, showcase that our model, UNIMO-G, consistently surpasses baseline models in handling both two and three-entity multi-modal prompts. This clear performance advantage underscores the effectiveness of our approach across varying levels of complexity.
@@ Appendix C Analysis of Multi-modal Prompts
Appendix D Images and Prompts of MultiBench